# StructLM: Towards Building Generalist Models for Structured Knowledge Grounding

### v2

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the CodeLlama architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models on 14 out of 18 evaluated datasets and establishes new SoTA achievements on 7 SKG tasks. Furthermore, StructLM demonstrates exceptional generalization across 6 novel SKG tasks. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level.
@@ 1 Introduction
Traditionally, users need to write programs to interface with structured data like tables, databases, knowledge graphs, etc. It requires the users to master the domain-specific language like SQL, SPARQL, etc. Recently, researchers have explored the possibility of automating the interface with natural language to enable potential use cases in question-answering Pasupat and Liang (2015); Zhong et al. (2017); Nan et al. (2022), summarization Parikh et al. (2020); Nan et al. (2021); Bao et al. (2018), and fact verification Aly et al. (2021); Chen et al. (2019); Gupta et al. (2020a), among others, all grounded to a structured knowledge source. This effort can lower the barrier for end users to access massive amount of structured data.
@@
Figure 1: StructLM can ground on structured and unstructured knowledge to respond to human queries. The previous SoTA was attained by many different task-specific models like TAPEX Liu et al. (2021), USKG Xie et al. (2022), TableLlama Zhang et al. (2023), BINDER-Codex Cheng et al. (2022), etc. StructLM (a single model) beats the previous SoTAs on seven out of eighteen SKG tasks.
@@
The previous work Yu et al. (2020); Liu et al. (2021); Xie et al. (2022); zhang2023tablellama has been mostly focused on building task-specific models for different tasks with rather limited generalization ability. Building a generalist structure knowledge grounding (SKG) system across a wide range of tasks proves to be challenging. This is mainly due to the heterogeneity of data format and use cases. We evaluated GPT-3.5-Turbo Jiang et al. (2023) on 18 SKG tasks and observed that its performance is on average 35% lower than the SoTA specialized models. It shows that the LLM’s ability on SKG is heavily overlooked during the pre-training phase.
@@
Figure 2: Overview of StructLM. This figure illustrates the prompting structure of StructLM, highlighting its capability to process various forms of structured data beyond linearized data tables, including linearized database schemas and knowledge graphs. StructLM is also assessed on held-out tasks that bear similarity to groups of held-in tasks, but also differences that must be overcome.
@@
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the CodeLlama family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin.
@@
  • We show that StructLM is able to show strong zero-shot generalization capability on unseen structure knowledge grounding tasks, which was not shown by previous models.
  • We find that scaling general instruction-tuning data improves generalization ability, and that code-pretrained base models indeed improve model performance on the SKG tasks.
2 Related Work
@@ 2.1 Solving SKG tasks
Structured knowledge, such as web tables, knowledge graphs, and databases, have long been the subject of study in knowledge grounding. However, SKG tasks have heterogeneous data formats which have inspired methods that leverage specific training setups to learn those representations. For example, PTab Liu et al. (2022) and MultiHiertt Zhao et al. (2022) learn the contextual representation of tabular data by incorporating semantic information through specific training methods or reasoning approaches. RASAT Qi et al. (2022a) integrates relation-aware self-attention with the Transformer seq2seq architecture and utilizes various relational structures to address SQL problems. TAPEX Liu et al. (2021) conducts pretraining over tabular/database data with the help of an SQL executor to provide supervision.
More recently, methods have begun to move away from these auxiliary task-specific structures. USKG Xie et al. (2022) were the first to unify many SKG tasks into a sequence-to-sequence format, allowing them to be aggregated into the same data mixture. StructGPT Jiang et al. (2023) represents a line of work that uses prompting frameworks on powerful LLMs to solve tasks with more robustness and accuracy. In contrast, our work examines open models and tries to assess their fundamental capabilities. Contemporary to our work, TableLlama Zhang et al. (2023) has argued that tabular data deserves special attention. Focusing on this domain, their method fine-tunes on several new tabular tasks to improve table understanding, and operates on a longer 8k context length. These improvements can be additive to our work.
USKG showed the benefit of the sequence-to-sequence format unification in more powerful language models, however, it was not able to show strong advantages to constructing a multi-task mix of SKG data over single-task models or task-specific training methods such as prefix-tuning in terms of performance. We seek to address this gap with our construction of an instruction-tuning SKG dataset with a large subset of USKG’s tasks, and the resulting model, StructLM, which does show strong multi-task performance.
@@ 2.2 LLMs with Instruction Tuning
Instruction-tuning (IT) has been popularized as a method to address the gap between training objectives and user goals in LLMs. This technique involves additional training of LLMs using pairs of instructions and outputs. IT enhances both the controllability and the predictability of the models, aligning them more closely with user expectations. Furthermore, recent studies such as FLAN Wei et al. (2022), UL2 Tay et al. (2023a), and Llama2 Touvron et al. (2023) have shown that IT can improve the performance of downstream tasks through multi-task learning across diverse data types. While FLAN-UL2 trains on a subset of 11 tasks from USKG, it also trains on many more unrelated language tasks. In our work, by focusing on SKG data, we hope to provide a focused study that can act as a reference for future work to improve performance on this task type.
@@ 2.3 Reasoning Capability in LLMs
Reasoning stands as a pivotal skill for LLMs in the development of real-world AI applications which would enable the autonomous completion of many thought-intensive tasks viewed traditionally to require human thinking, like programming or mathematical problem-solving Li et al. (2022). Recent studies Li et al. (2022, 2023c); Rozière et al. (2023); Azerbayev et al. (2023a) indicate that LLMs trained on code and mathematical datasets exhibit profound reasoning skills, and can even achieve performance on par with human levels. For example, CodeLlama Rozière et al. (2023), a foundation model trained on more programming data, has significantly improved reasoning capabilities across a variety of programming and mathematical benchmarks. Furthermore, Llemma Azerbayev et al. (2023a) continues to pretrain the CodeLlama model on a mix of scientific papers, math-related web data, and mathematical code. Its results show excellent reasoning capabilities on the MATH benchmark Hendrycks et al. (2021) and the ability to prove theorems without further fine-tuning. On the fine-tuning side, WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) have shown the effectiveness of instruction tuning on reasoning capabilities, given high quality data.
@@
Figure 3: Breakdown of Structured Knowledge Types and Tasks. On the inside, we see a coarse breakdown of the different categories of structured inputs in our dataset. On the outside, we see the tasks that are represented for those structured knowledge types. We also have a considerable proportion of general instruction following data from SlimOrca Lian et al. (2023).
@@ 3.1 Dataset Curation
Motivated by the goal of training a language model (LM) generally capable of a wide range of structured data tasks, we select a total of $25$ SKG tasks to study. We report results on $18$ held-in and $6$ held-out tasks, where each held-out task meant to roughly evaluate the generalization capability of a held-in task group. In total, our held-in training dataset contains approximately $700$k SKG examples. We describe the held-in dataset groups below.
@@
Knowledge-grounded Conversations. This group of tasks evaluates knowledge grounding in-conversation. Humans naturally interface with LMs is through chat, and enabling this capability can lower the barrier to accessing the information in stored structured data. These tasks track user intention through provided dialogue and ask the model to provide an answer to the latest question. The held-out dataset for this task group is CoSQL.
@@
SQL or domain-specific languages SQL is the language most commonly used to interface with structured data today. Understanding how to write SQL also requires understanding of abstractions of tables and how they are linked together. In other domain-specific languages, the MTOP task measures a model’s ability to parse a specification and generate an API call, which sees potential in LLM tool use (e.g., Qin et al. (2023)). The corresponding held-out dataset for this task group is intended to be BIRD Li et al. (2023b), which further tests SQL generation abilities.
@@
General instruction data. In addition to the SKG datasets within the held-in dataset mixture, we also included general instruction tuning data without any structured knowledge component, to maintain the instruction-following ability of our model. We use SlimOrca Lian et al. (2023), which is constructed from cleaned GPT-4 responses to a number of prompts from existing general large-scale instruction-following datasets. We detect no signs of data contamination for our held-out datasets based on our ablation results. We give a detailed overview of all dataset statistics in  Table 1.
@@ 3.2 Instruction Finetuning Approach
To instruction tune our model, each example in our dataset consists of a system prompt, instruction, input, and output. For all SKG data examples, we use the same system prompt. For each dataset, we write 10 instruction variations, which are randomized when constructing the training samples. For SKG data, the input is composed of a combination of a structured knowledge input and accompanying text that could be a question, statement, or anything that would be required to specify the task. The exact prompt format is provided in Figure 6.
@@ 3.3 Training and Evaluation Details
The base models for StructLM are the CodeLlama-Instruct family of models Rozière et al. (2023). We finetune all models with a batch size of $512$ for $3$ epochs on A800 gpus. This training setup is largely in line with community conventions, such as the settings used for the WizardLM Xu et al. (2023), WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) models.
@@
We follow the structured data linearization conventions in USKG Xie et al. (2022). However, we use a different truncation scheme as described below. During training, we maintain a maximum sequence length of $2048$. USKG notes that training and running inference with longer context is beneficial. As such, when truncating, we consider the combined token length of the prompt input and output label. We truncate only the structured knowledge portion of the input so that the example becomes at most $2048$ tokens long. As shown in the dataset statistics in Table 1, setting the max token length of the examples in our dataset to $2048$ allows nearly all examples to fit within the context window with rare truncations. We discard examples for which even this structured input truncation is insufficient (e.g. the output is too long). During inference, we set the input token length to $2048$, to allow even more structured information to be placed within the input context. We set the maximum generation length to 1024, which is sufficient for all correct responses in all datasets. For each model, including our single-task finetuned models, we choose the best performing checkpoint of the $3$ checkpoints produced at the end of each epoch.
@@ 4 Experiments
ToTTo BLEU 49.9 20.7 17.5 48.8 - - 49.0 49.4 49.3 50.2 +1.2
GrailQA EM 77.1 9.3 1.0 77.0 - - 70.1 80.4 79.2 82.2 +12.1
@@
Spider EM 80.5 43.8 5.2 67.3 - - 71.8 72.4 74.1 74.6 +2.8
KVRet All Micro 67.9 52.9 39.5 70.9 - - 67.9 72.6 69.5 69.3 +4.5
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 - 59.4 59.2 59.1 61.1 +1.7
SParC EM 68.2 32.2 3.2 62.3 - - 61.5 61.9 64.9 63.4 +3.4
CompWebQ Acc 76.8 48.9 3.1 75.6 75.9 - 73.3 78.3 80.4 81.9 +8.6
TabFact Acc 93.0 62.4 0.0 79.6 87.1 82.5 83.7 80.8 84.7 86.6 +2.9
WikiTQ All Ex 65.9 24.8 0.2 45.7 54.6 - 49.3 50.1 53.4 55.7 +6.4
WikiSQL All Ex 93.0 31.5 0.4 86.5 87.3 - 86.0 88.7 87.2 87.6 +1.7
FeTaQA BLEU 39.0 7.4 5.6 33.8 35.8 39.0 33.4 36.0 35.6 37.5 +4.1
FEVEROUS Acc 85.6 57.8 58.4 78.1 85.6 - 82.4 84.4 85.0 85.7 +3.3
MultiWOZ Joint Acc 60.6 8.9 0.0 53.0 - - 55.4 54.5 53.0 53.8 -0.9
DART BLEU 52.0 59.0 54.6 60.3 50.4 - 46.7 62.2 61.4 61.8 +15.5
Logic2Text Blec 95.3 78.5 59.1 89.5 - - 91.4 88.9 90.1 89.1 -1.3
MTOP EM 87.5 1.4 0.0 77.4 87.5 - 86.8 81.2 81.6 82.1 -4.7
Average 74.9 39.5 20.2 68.2 - - 69.3 71.1 71.3 72.6 +3.3
@@
BIRD Acc 36.6* 21.8 0.0 24.4* - - - 22.3 22.8 24.7 -
CoSQL EM 58.3* 33.7 0.2 52.4* - - - 49.8 52.2 55.0 -
SQA Acc 70.5* 18.7 2.3 60.4* 70.1* - - 49.7 36.1 44.2 -
Infotabs Acc 75.6* 46.9 40.2 68.7* - 35.5 - 55.3 58.1 61.8 -
WikiTableText BLEU 33.7* 3.8 5.7 39.8* - 10.2 - 8.3 9.3 8.8 -
Finqa Acc 71.1* 31.4 1.7 79.7* - 18.6 - 27.3 25.6 36.2 -
Average 57.6* 26.1 8.4 54.2* - - - 35.5 34.0 38.4 -
Table 2: The overall evaluation results of our model against other baselines. Cells with "-" in the held-in part mean that the model was trained on these datasets, thus the results are not comparable. Cells with "-" in the held-out part mean the results are not yet available. Cells in the held-out section with "*" are held-in results and other than ST, are copied from the original papers for reference. TableLlama held-out results are obtained by running their checkpoint. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B. ST refers to the single-task finetuning result of CodeLlama-Instruct-7B on each dataset. XYZ denotes the state-of-the-art score achieved by us. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
To evaluate the benefits of our instruction-tuning dataset mix, we also run single-task baseline (each a 7B model) on each task and report their individual performance. We again use CodeLlama-7B-Instruct as the base model for each, and match each single task model on the same number of epochs (3) that was used to train the multitask models, ensuring that each model has seen the same data the same number of times. We observe that our multi-task models outperform these single-task models on nearly every task, with some by a considerable margin of up to $7\%$. This demonstrates the effectiveness of our instruction tuning dataset and supports the presence of cross-task generalization within our dataset mix.
When compared to the single-task 18 USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although the single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B Tay et al. (2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Our results on held-in datasets (Tabfact and FeTaQA) are on par with TableLlama Zhang et al. (2023), which is an LLM pre-trained on 2.6M table understanding tasks.
@@
On held out tasks, StructLM shows strong generalization performance, outperforming ChatGPT on 5 of 6 tasks. We note that our zero-shot prompting method for the FinQA Chen et al. (2021) dataset requires the model to generate custom code and is focused only on the financial domain, and the input to Infotabs Gupta et al. (2020a) is a different type of table structure unseen in the input data, with different labels for the boolean output categories than the choices from the held-in fact verification tasks. This generalization capability is non-existent in USKG models as the name of each held-in task is specified in their training inputs.

### v3

@@ Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the CodeLlama architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models (Xie et al., 2022) on 14 out of 18 evaluated datasets and establishes new SoTA performance on 7 SKG tasks. Furthermore, StructLM demonstrates strong generalization across 6 novel held-out SKG tasks, outperforming TableLlama by an average of 35% and Flan-UL2 20B by an avarage of 10%. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level. We release the model weights and training dataset to the community, along with relevant code on Github.
@@ 1 Introduction
Traditionally, users need to write programs to interface with structured data like tables, databases, knowledge graphs, etc. This often requires that they master a the domain-specific language like SQL, SPARQL, etc. or develop domain specific skills for that structured type. Recently, researchers have explored the possibility of automating the interface with natural language to enable potential use cases in question-answering (Pasupat & Liang, 2015; Zhong et al., 2017; Nan et al., 2022), summarization (Parikh et al., 2020; Nan et al., 2021; Bao et al., 2018), and fact verification (Aly et al., 2021; Chen et al., 2019; Gupta et al., 2020b), among others, all grounded to a structured knowledge source. This effort can lower the barrier for end users to access a massive amount of structured data.
@@
Figure 1: StructLM can ground on structured and unstructured knowledge to respond to human queries. The previous SoTA was attained by many different task-specific models like TAPEX (Liu et al., 2021), USKG (Xie et al., 2022), TableLlama (Zhang et al., 2023), BINDER-Codex (Cheng et al., 2022), etc. StructLM beats the SoTAs on seven SKG tasks.
@@
Previous work (Yu et al., 2020; Liu et al., 2021; Xie et al., 2022; Zhang et al., 2023) has been mostly focused on building task-specific models for different tasks with rather limited generalization ability. Building a generalist structure knowledge grounding (SKG) system across a wide range of tasks proves to be challenging. This is mainly due to the heterogeneity of data format and use cases. We evaluated GPT-3.5-Turbo (Jiang et al., 2023) on 18 SKG tasks and observed that its performance is on average 35% lower than the SoTA specialized models. It shows that the LLM’s ability on SKG is heavily overlooked during pre-training.
@@
Figure 2: Overview of StructLM. This figure illustrates the prompting structure of StructLM, highlighting its capability to process various forms of structured data beyond linearized data tables, including linearized database schemas and knowledge graphs.
@@
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the CodeLlama family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin. Morever, we show that compared with TableLlama, a recent open language model finetuned for tabular inference tasks, our held-out performance is vastly superior. Compared with Flan-UL2-20B, our models also generalize better overall on the held out tasks by as much as 10% on average.
@@
  • We show that StructLM is able to show strong zero-shot generalization capability on unseen structure knowledge grounding tasks, which was not shown by previous models. StructLM outperforms Flan-UL2 20B on 3 of 5 tasks and TableLlama on all 6 tasks. The average absolute improvement over them are 10% and 35%.
  • We find that mixing in general-domain instruction-tuning data during finetuning preserves generalization ability, and that code-pretrained base models can improve model performance on the SKG tasks.
2 Related Works
@@ 2.1 Solving SKG tasks
Structured knowledge, such as web tables, knowledge graphs, and databases, have long been the subject of study in knowledge grounding. However, SKG tasks have heterogeneous data formats which have inspired methods that leverage specific training setups to learn those representations. For example, PTab (Liu et al., 2022) and MultiHiertt (Zhao et al., 2022) learn the contextual representation of tabular data by incorporating semantic information through specific training methods or reasoning approaches. RASAT (Qi et al., 2022a) integrates relation-aware self-attention with the Transformer seq2seq architecture and utilizes various relational structures to address SQL problems. TAPEX (Liu et al., 2021) conducts pretraining over tabular/database data with the help of an SQL executor to provide supervision.
More recently, methods have begun to move away from these auxiliary task-specific structures. USKG (Xie et al., 2022) were the first to unify many SKG tasks into a sequence-to-sequence format, allowing them to be aggregated into the same data mixture. However, they were not able to show strong performance improvements to constructing a multi-task mix of SKG data over task-specific tuning methods. StructGPT (Jiang et al., 2023) represents a line of work that uses prompting frameworks on powerful LLMs to solve tasks with more robustness and accuracy. In contrast, our work examines open models and tries to assess their fundamental capabilities. Contemporary to our work, TableLlama (Zhang et al., 2023) has argued that tabular data deserves special attention. Focusing on this domain, their method fine-tunes on several new tabular tasks to improve table understanding, and operates on a longer 8k context length. These improvements can be additive to our work.
@@ 2.2 LLMs with Instruction Tuning
Instruction-tuning (IT) has been popularized as a method to address the gap between training objectives and user goals in LLMs. This technique involves additional training of LLMs using pairs of instructions and outputs. IT enhances both the controllability and the predictability of the models, aligning them more closely with user expectations. Furthermore, recent studies such as FLAN (Wei et al., 2022), UL2 (Tay et al., 2023a), and Llama2 (Touvron et al., 2023) have shown that IT can improve the performance of downstream tasks through multi-task learning across diverse data types. While FLAN-UL2 trains on a subset of 11 SKG tasks, it also trains on many more unrelated language tasks. In our work, by focusing on SKG data, we hope to provide a focused study that can act as a reference for future work to improve performance on this task type.
@@ 2.3 Reasoning Capability in LLMs
Reasoning stands as a pivotal skill for LLMs in the development of real-world AI applications which would enable the autonomous completion of many thought-intensive tasks viewed traditionally to require human thinking, like programming or mathematical problem-solving (Li et al., 2022). Recent studies (Li et al., 2022; 2023c; Rozière et al., 2023; Azerbayev et al., 2023a) indicate that LLMs trained on code and mathematical datasets exhibit profound reasoning skills, and can even achieve performance on par with human levels. For example, CodeLlama (Rozière et al., 2023), a foundation model trained on more programming data, has significantly improved reasoning capabilities across a variety of programming and mathematical benchmarks. Furthermore, Llemma (Azerbayev et al., 2023a) continues to pretrain the CodeLlama model on a mix of scientific papers, math-related web data, and mathematical code. Its results show excellent reasoning capabilities on the MATH benchmark (Hendrycks et al., 2021) and the ability to prove theorems without further fine-tuning. On the fine-tuning side, WizardMath (Luo et al., 2023a), and WizardCoder (Luo et al., 2023c) have shown the effectiveness of instruction tuning on reasoning capabilities, given high quality data.
@@
Figure 3: Breakdown of Structured Knowledge Types and Tasks in the Training Data. On the left side, we see a coarse breakdown of the different categories of structured inputs in our dataset. On the right side, we see an overview of the task groups that are represented for those structured knowledge types.
@@ 3.1 Dataset Curation
Motivated by the goal of training a language model generally capable of a wide range of structured data tasks, we select a total of $25$ SKG tasks to study. We report results on $18$ held-in and $6$ held-out tasks, where each held-out task meant to roughly evaluate the generalization capability of a held-in task group. In total, our held-in training dataset contains approximately $700$k SKG examples. We describe the held-in dataset groups below.
@@
Knowledge-grounded Conversations. This group of tasks evaluates knowledge grounding in-conversation. Humans naturally interface with LMs through chat, and enabling this capability can lower the barrier to accessing the information in stored structured data. These tasks track user intention through provided dialogue and ask the model to provide an answer to the latest question. The held-out dataset for this task group is CoSQL.
@@
SQL or domain-specific languages SQL is the language most commonly used to interface with structured data today. Understanding how to write SQL also requires understanding of abstractions of tables and how they are linked together. In other domain-specific languages, the MTOP task measures a model’s ability to parse a specification and generate an API call, which sees potential in LLM tool use (e.g., (Qin et al., 2023)). The corresponding held-out dataset for this task group is intended to be BIRD (Li et al., 2023b), which further tests SQL generation abilities.
@@
General instruction data. In addition to the SKG datasets within the held-in dataset mixture, we also included general instruction tuning data without any structured knowledge component, to maintain the instruction-following ability of our model. We use SlimOrca (Lian et al., 2023), which is constructed from cleaned GPT-4 responses to a number of prompts from existing general large-scale instruction-following datasets. We detect no signs of data contamination for our held-out datasets based on our ablation results. We give a detailed overview of all dataset statistics in  Table 1.
@@ 3.2 Instruction Finetuning Approach
To instruction tune our model, each example in our dataset consists of a system prompt, instruction, input, and output. For all SKG data examples, we use the same system prompt. For each dataset, we write 10 instruction variations, which are randomized when constructing the training samples. For SKG data, the input is composed of a combination of a structured knowledge input and accompanying text that could be a question, statement, or anything that would be required to specify the task. The prompt is provided in Figure 6.
@@ 3.3 Training and Evaluation Details
The base models for StructLM are the CodeLlama-Instruct family of models (Rozière et al., 2023). We finetune all models with a batch size of $512$ for $3$ epochs on A800 gpus. We train our 7-34B models on 16-64 GPUs using DeepSpeed ZeRO-3 (Rasley et al., 2020) such that the total training time for each model is between 3-5 days. This training setup is largely in line with community conventions, such as the settings used for the WizardLM (Xu et al., 2023), WizardMath (Luo et al., 2023a), and WizardCoder (Luo et al., 2023c) models.
@@
We follow the structured data linearization conventions in USKG (Xie et al., 2022). However, we use a different truncation scheme as described below. During training, we maintain a maximum sequence length of $2048$. To preserve as much input context as possible, when truncating we consider the combined token length of the prompt input and output label. We truncate only the structured knowledge portion of the input so that the example becomes at most $2048$ tokens long. As shown in the dataset statistics in Table 1, setting the max token length of the examples in our dataset to $2048$ allows nearly all examples to fit within the context window with rare truncations. We discard examples for which even this structured input truncation is insufficient (e.g. the output is too long). During inference, we set the input token length to $2048$, to allow even more structured information to be placed within the input context. We set the maximum generation length to 1024, which is sufficient for all correct responses in all datasets. For each model, including our single-task finetuned models, we choose the best performing checkpoint of the $3$-epoch checkpoints.
@@ 4 Experiments
ToTTo BLEU 49.9 20.7 17.5 48.8 - - 49.0 49.4 49.3 50.2 +0.3
GrailQA EM 77.1 9.3 1.0 77.0 - - 70.1 80.4 79.2 82.2 +5.1
@@
Spider EM 80.5 43.8 5.2 67.3 - - 71.8 72.4 74.1 74.6 -5.9
KVRet All Micro 67.9 52.9 39.5 70.9 - - 67.9 72.6 69.5 69.3 +4.7
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 - 59.4 59.2 59.1 61.1 -7.3
SParC EM 68.2 32.2 3.2 62.3 - - 61.5 61.9 64.9 63.4 -3.3
CompWebQ Acc 76.8 48.9 3.1 75.6 75.9 - 73.3 78.3 80.4 81.9 -5.1
TabFact Acc 93.0 62.4 0.0 79.6 87.1 82.5 83.7 80.8 84.7 86.6 -6.4
WikiTQ All Ex 65.9 24.8 0.2 45.7 54.6 - 49.3 50.1 53.4 55.7 -10.2
WikiSQL All Ex 93.0 31.5 0.4 86.5 87.3 - 86.0 88.7 87.2 87.6 -5.4
FeTaQA BLEU 39.0 7.4 5.6 33.8 35.8 39.0 33.4 36.0 35.6 37.5 -1.5
FEVEROUS Acc 85.6 57.8 58.4 78.1 85.6 - 82.4 84.4 85.0 85.7 +0.1
MultiWOZ Joint Acc 60.6 8.9 0.0 53.0 - - 55.4 54.5 53.0 53.8 -6.1
DART BLEU 52.0 59.0 54.6 60.3 50.4 - 46.7 62.2 61.4 61.8 +10.2
Logic2Text Blec 95.3 78.5 59.1 89.5 - - 91.4 88.9 90.1 89.1 -5.2
MTOP EM 87.5 1.4 0.0 77.4 87.5 - 86.8 81.2 81.6 82.1 -5.4
Average 74.9 39.5 20.2 68.2 - - 69.3 71.1 71.3 72.6 -2.2
@@
BIRD Acc 36.6* 21.8 0.0 24.4* 1.0 0 0 22.3 22.8 24.7 +2.9
CoSQL EM 58.3* 33.7 0.2 52.4* 5.1 0 0 49.8 52.2 55.0 +21.3
SQA Acc 70.5* 18.7 2.3 60.4* 70.1* 0 0 49.7 36.1 44.2 +31
Infotabs Acc 75.6* 46.9 40.2 68.7* 70.3 16.2 0 55.3 58.1 61.8 -7.1
WikiTableText BLEU 33.7* 3.8 5.7 39.8* 19.4 3.4 0 8.3 9.3 8.8 -9.9
Finqa Acc 71.1* 31.4 1.7 79.7* 5.9 2.6 0 27.3 25.6 36.2 +4.8
Average 57.6* 26.1 8.4 54.2* 28.6* 3.7 0 35.5 34.0 38.4 +10.3
Table 2: The overall evaluation results of our model against other baselines. Cells with ”-” in the held-in part mean that the model did not train on this dataset, and results are not comparable. USKG models are overfit to the held-in dataset labels, and thus cannot generalize comparably. Cells in the held-out section with ”*” are held-in results. SoTA results are copied from the original papers for reference. ST refers to the single-task fine-tuning result of CodeLlama-Instruct-7B on each dataset. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B, with the same structure input format as StructLM. $\Delta$ refers to the difference between StructLM and the best known result. score denotes the state-of-the-art score on specific tasks. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
@@
To evaluate the benefits of our instruction-tuning dataset mix, we also run single-task baseline (each a 7B model) on each task and report their individual performance. We again use CodeLlama-7B-Instruct as the base model for each, and match each single task model on the same number of epochs (3) that was used to train the multitask models, ensuring that each model has seen the same data the same number of times. We observe that our multi-task models outperform these single-task models on nearly every task, with some by a considerable margin of up to $7\%$. This demonstrates the effectiveness of our instruction tuning dataset and supports the presence of cross-task generalization.
When compared to the 18 task-specific USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although the single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B (Tay et al., 2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Our results on held-in datasets (Tabfact and FeTaQA) are on par with TableLlama (Zhang et al., 2023), which is an LLM pre-trained on 2.6M table understanding tasks. We are vastly outperforming TableLlama on the held-out datasets.
@@
On held out tasks, StructLM shows strong generalization performance, outperforming ChatGPT on 5 of 6 tasks. These novel tasks contain remarkably different data format than the training dataset. For example, FinQA (Chen et al., 2021) requires the model to generate a mathematical expression to answer a question in the financial domain, while BIRD (Li et al., 2023b) requires the model to interface with multiple databases. These skills are not exactly covered in the training set, thus they put high demand for the models’ generalization capabilities. Such generalization capabilities are non-existent in USKG models as they are task-specific. Another table foundation model TableLlama (Zhang et al., 2023) would also fail on most of the held-out tasks with performance close to zero. In contrast, StructLM is much more capable of generalization on these novel tasks. The average improvement over TableLlama is 35%. Compared to another strong foundation structure-knowledge-grounding model Flan-UL2 20B (Tay et al., 2023b), we also outperform it on held-out tasks by an average of 10%.